OpenStudy (yrelhan4):

a double slit is illuminated by two wavelengths 450nm and 600nm. What is the lowest order at which the maxima of one wavelength coincides with minima of the other? what i did is n*450=[n-(1/2)]*600 this gives me 2. answer is 3. whats wrong?

OpenStudy (yrelhan4):

@Mashy

OpenStudy (yrelhan4):

@shubhamsrg

OpenStudy (shubhamsrg):

well your method seems right to me.

OpenStudy (yrelhan4):

it looks right to me too. i checked it on google. got a couple of google books results which have the answer 3 too. must be something wrong with my logic.

OpenStudy (shubhamsrg):

Leme try again.

OpenStudy (shubhamsrg):

wait, what was I saying,nopes the method is wrong. :P

OpenStudy (yrelhan4):

hmm. a book has a similar question. so i followed that. :P

OpenStudy (shubhamsrg):

maxima for 450 = m(450) minima for 450 = (m -1/2)450 maxima for 600 = n(600) minima for 600 = (n- 1/2)600 thus, either 450m = (n- 1/2)600 or 600n = (m- 1/2)(450) => m/(n -1/2) = 4/3 and n/(m- 1/2) = 3/4 on comparison,n and m are symmetrical . hit and trial shows n=2 and m=2 satisfy it ans = 4 lol, my method is thus flawed, sorry about that.

OpenStudy (yrelhan4):

nice try. better than mine for sure.

OpenStudy (shubhamsrg):

I did what you have done. But, you can not take both orders to be n, your eqn is correct with one n and other m. Order then = n+m, but but, it fails! :3

OpenStudy (yrelhan4):

hmm.

OpenStudy (shubhamsrg):

can you tell me when is minima = (n -1/2)lambda and when is it (n +1/2)lambda ?

OpenStudy (shubhamsrg):

Reason I ask is (n+ 1/2) guides us to the ans i.e. 3! :P

OpenStudy (yrelhan4):

i have always used n-1/2. n+1/2 is used for diffraction max. i can be wrong.

OpenStudy (shubhamsrg):

can't * ;)

OpenStudy (yrelhan4):

haha. if that yields the answer i am obviously wrong. :P

OpenStudy (shubhamsrg):

as according to him, I'll take n+1/2 and thus, 3 is the ans! :P

OpenStudy (shubhamsrg):

But am not too comfortable with it,

OpenStudy (shubhamsrg):

@Mashy @Vincent-Lyon.Fr @DLS (the champ)

OpenStudy (yrelhan4):

i dont know. i checked in 2 books now. they have n-1/2. and champ is still in 11th. :P

OpenStudy (shubhamsrg):

hmm,must be some problem in the logic then.

OpenStudy (yrelhan4):

yeah. lets see.

OpenStudy (anonymous):

its n + 1/2 lamda.. when n starts from 0,1,2,3 etc its n -1/2 lambda when n starts from 1, 2,3 etc.. you can take anything however.. since in this case we are considering both min and max together, for max we usually start with 0, i prefer to take the formula which allows me to start with 0 for min too.. so i would go with n+1/2

OpenStudy (anonymous):

so when i do 450m = 600 (n +1/2) i get m = 2 and n = 1 which means second order maxima of 450nm light coincides with 2nd order minima (remember we have chosen 0 as the first order minima) of the 600nm light.. i dunno how shubhang got 3? :O

OpenStudy (shubhamsrg):

well n and m are respective orders right? n=1 and m=2 , so net =3 ?

OpenStudy (shubhamsrg):

Well I may be wrong, I am not too comfortable with wave optics at the moment.

OpenStudy (shubhamsrg):

shubhang? o.O

OpenStudy (yrelhan4):

so is this net thing right?

OpenStudy (anonymous):

what net thing?? .. the question is quite ambigious :-/.. the question doesn't really say we need to add up the orders :-/

OpenStudy (anonymous):

and besides i got.. both 2nd order .. ! :P

OpenStudy (shubhamsrg):

That is what I had yielded in the first place! B|

OpenStudy (yrelhan4):

think about it.. you usually get the answer after some time. take your time. :P also if we dont take n,m and just do n*600=(n+1/2)*450, we get n=2. i know its not correct but i have done this in questions involving nth fringe where we do n*a=(n+1)*b where nth fringe coincides. :P

OpenStudy (yrelhan4):

where a and b are fringe widths.

OpenStudy (anonymous):

well thats because.. it turns out the second orders DO CONINCIDE.. if the second order's didn't conincide.. then in your equation you would get n as some fraction!

OpenStudy (yrelhan4):

hmmmm.. got that part. and what do you think of n=3?

OpenStudy (yrelhan4):

look at this question. why have they taken n only?

OpenStudy (shubhamsrg):

I will not tell you.

OpenStudy (yrelhan4):

:O was that a stupid question?

OpenStudy (anonymous):

well thats because the incident light itself is made up of two colors you ll see a maxima on the screen, when the maximas of BOTH the yellow light conincide .. !

OpenStudy (shubhamsrg):

nopes, because I don't know! -_-

OpenStudy (anonymous):

in this quesiton.. we have two seperate patterns itself..

OpenStudy (yrelhan4):

hmm i see.

OpenStudy (anonymous):

wait.. what?.. maxima co inciding with minima!!.. wait wait wait!

OpenStudy (shubhamsrg):

take your time! :P

OpenStudy (anonymous):

they have mentioned.. that until the maxima of one conincides with minima of the other, you ll see pattern.. after that you won't.. quite frankly i dunno why? but atleast it makes more sense then what your book says!

OpenStudy (anonymous):

oh wait.. they both say the same thing.. nv :D

OpenStudy (yrelhan4):

lol. have some water sir.

OpenStudy (shubhamsrg):

I really got to revise my wave optics! hmm.. :(

OpenStudy (anonymous):

this IIT physics is so stupid!.. its basically puzzles.. who wants to solve puzzles.. :-/

OpenStudy (shubhamsrg):

omg! look what I found http://inst.eecs.berkeley.edu/~ee119/sp10/Homework/HW11SOL.pdf

OpenStudy (shubhamsrg):

see "4. Young's double slit experiment" part d)

OpenStudy (shubhamsrg):

and leme bring to everyone;s notice, the CHAMP has arrived! \m/

OpenStudy (dls):

im getting 3 :/

OpenStudy (shubhamsrg):

I am not surprised !

OpenStudy (anonymous):

but m2 = 1 is second order minima :( :( .. not first order..

OpenStudy (dls):

wait a sec im posting my solution

OpenStudy (yrelhan4):

NOW THATS WHY YOU ARE A CHAMP. explain, NOW!

OpenStudy (shubhamsrg):

@Mashy :(

OpenStudy (dls):

Enjoy,goodbye.

OpenStudy (yrelhan4):

WTF. :/

OpenStudy (dls):

maybe i can help next year..:P

OpenStudy (dls):

answer is 2.

OpenStudy (dls):

you have a misprint

OpenStudy (shubhamsrg):

we have a new answer then! B|

OpenStudy (yrelhan4):

hmm. i guess its a misprint then. so the link shubham posted was also wrong?

OpenStudy (dls):

are you talking about yahoo answers?

OpenStudy (anonymous):

MUHAHAHAHHAA see i was right all along!!

OpenStudy (anonymous):

mashy mashy he is the man if he can't do it nobody can :D

OpenStudy (yrelhan4):

what if they say first order minima? whats n?

OpenStudy (anonymous):

see first order minima means first minima.. thats when the path difference is exactly equal to lambda /2 so you see.. that can happen if you plug in n=1 in the equation n-1/2 lambda.. or plug in n=0 in equation n+1/2 lamda.. so it really depends on what equation you choose.. but ultimately you see its the same damn thing :P

OpenStudy (yrelhan4):

i got it, i guess. you are telling me to see that if i use n-1/2, then i see, first order minima is n=1. and if, i see, we use n+1/2, then first order minima is n=0. i see its the same damn thing.

OpenStudy (anonymous):

i see that you are trying to be funny :P.. !! now lemme take bath !

OpenStudy (yrelhan4):

not really. :P

OpenStudy (shubhamsrg):

question no.55 answer is c)

OpenStudy (shubhamsrg):

LAWL

OpenStudy (yrelhan4):

oh dang. whatever now. must be a misprint. :P

OpenStudy (shubhamsrg):

well 2 of the books say 3, one says 2 .. so you all can calculate the respective probabilities! ;)

OpenStudy (yrelhan4):

i'll go with the logic. :P and if the same question comes in the exam, i'll probably leave it. :P

OpenStudy (shubhamsrg):

hey, how did "LAWL" come down? o.O

OpenStudy (yrelhan4):

hmm. good question.

OpenStudy (shubhamsrg):

I'll revise my wave optics and re-visit the question in the near future.

OpenStudy (shubhamsrg):

Though I don't guarantee I'll be able to resolve the conflict! -_-

OpenStudy (yrelhan4):

haha. bye.